# Make Me Happier: Evoking Emotions Through Image Diffusion Models

### v1

@@ 4 Experiment
We selected 504 images randomly from the EmoSet dataset [31], ensuring they were distinct from those in our EmoPair dataset. According to the emotion labels of EmoSet, there are 63 images for each source emotion category. Subsequently, we evaluated both competitive baselines and our EmoEditor on these source images and presented the results in the following subsections.
@@
Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions. Considering emotional valence, we divided the test set into two subsets: within-valence emotion transfer and cross-valence emotion transfer. This division serves two purposes. Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween. Secondly, images of the same valence may contain subtle differences challenging for human viewers to discern. For instance, an amusement park scene might evoke a mixture of excitement and amusement.
@@
Baselines. We compare our EmoEditor with five state-of-the-art methods: (1) Color-transfer[21] (CT); (2) Neural-Style-Transfer[7] (NST). Both methods require an additional reference image from the target emotion category as inputs. We randomly selected these reference images from the EmoSet dataset. We also include representative text-to-image editing models (3) CLIP-Styler[14] (Csty) and (4) Ip2p [3]. We provide these models with the target emotion category as a one-word text prompt, such as “excitement". Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS). This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions. See Supp. for details.
Human Psychophysics Experiment. We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)[26]. We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. All participants are properly compensated. In each trial, participants engage in a “two-alternative forced choice" task. They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them. The two image stimuli consists of a result generated by our EmoEditor and a randomly sampled result from the five baselines introduced above. The trial presentation order and the binary choices for each trial are randomized. To assess the quality of data collection, we introduce control trials and use them as filtering criteria. We discard the data from those people who fail the control trials at least 17% of the time. See Supp. for more experimental details.
Evaluation Metrics. We introduce four novel evaluation metrics to quantitatively assess the emotion-evoked image generation performance of all methods. Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions. EMR: We use the emotion predictor $\mathcal{P}$ to predict the emotional category of generated images and calculate the proportion of alignment with the target emotions. ESR: Our emotion predictor $\mathcal{P}$ produces emotion probability distributions $e_{g}$ on generated images and $e_{s}$ on source images. We introduce Kullback-Leibler Divergence distance $KLD(\cdot)$ and calculate ESR as the proportion of the generated images, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$. $e_{oh}$ is the binary one-hot target emotion vector. It indicates the effect of edits introduced on the generated images in transitioning from the source emotion to the target emotion.
Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) measure the structural coherence and semantic consistency between source and generated images. ENRD: We first identify emotionally neutral regions on the source images by employing the Grad-CAM [25] technique through our emotion predictor $\mathcal{P}$. We then binarize the Grad-CAM maps with a threshold of 0.5, where a value of 1 on the resulting binary map signifies regions triggering emotions. By excluding these regions, we compute pixel-level L1 distance between emotionally neutral regions of the source and generated images. ESS: We apply the Canny edge detection algorithm on both source and generated images using thresholds of 200 and 500. Next, we compute the L1 norm between these edge maps to quantify their structural differences.
@@
[Uncaptioned image] Figure 5: The Human Psychophysics Experiment Results. The average proportion of images that human participants prefer our EmoEditor over other competitive methods is 56%. Chance is 50% in the red dotted line.    Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 32.61 7.36 NST 34.42 92.01 34.42 18.57 Csty 11.51 85.52 41.59 36.64 Ip2p 2.53 67.76 9.82 12.71 LMS 11.51 77.38 26.67 19.74 w/o $\mathcal{P}$ 5.06 69.15 20.08 14.93 w/o $L_{emb}$ 41.12 92.36 23.02 16.06 Ours 50.20 92.86 24.73 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor. See Sec. 4 for the methods and evaluation metrics. See Sec. 5.1 for the result analysis.
@@ 5.1 Quantitative Evaluation in Cross-Valence Scenarios
Figure 5 shows human psychophysics experimental results. If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks. Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec. 4. From Figure 5, we observed that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines. This preference indicates EmoEditor’s excellence in producing images that successfully evoke emotions aligning with the target emotions. This observation is also validated by the highest EMR and ESR scores of our EmoEditor in Table 1. Moreover, our EmoEditor achieves moderately lower ENRD and ESS, suggesting that the generated images by our EmoEditor also preserve structural similarity and semantic coherence.
From Figure 5, we also noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than just global adjustments to color and brightness. Indeed, CT also has lower EMR and ESR than our EmoEditor in Table 1. As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores. Similarly, our EmoEditor is significantly favored over Ip2p by human participants, indicating that current text-to-image diffusion models face challenges in understanding emotional cues and generating effective text instructions for editing emotion-evoked images. Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions.
Additionally, we noted that our EmoEditor also outperforms Large Model Series (LMS) by 4% based on human preference scores in Figure 5, 20% in EMR, 15% in ESR, 2 in ENRD and 3 in ESS in Table 1. It suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
Interestingly, we observed Neural-Style-Transfer [7] and CLIP-Styler [14] excel in inducing negative emotions through distortion and irregular textures on source images. Indeed, they are preferred by humans in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.). However, while these methods create a negative emotional atmosphere, they render the image challenging to comprehend. This is supported by the higher ENRD and ESS scores in Table 1, suggesting that the distorted images no longer preserve the structural similarities and semantic content of source images. In contrast, our EmoEditor outperforms both methods in generating positively valenced images, as supported by higher human preference scores in the lower quadrants of the confusion matrix (see Supp.).
In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance. This leads to lower EMR and ESR scores and higher ENRD and ESS scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions to modify source images, and executing iterative refinements during inference.
@@
Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Sec. 5.2 for the result analysis. See Supp. for more examples.
@@ 5.2 Visualization of Emotional Image Generation across Valence
Figure 6 presents visualizations of generated images for all competitive methods in cross-valence scenarios. Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation. For instance, in Row 3, Column 2, it enhances brightness but fails to evoke contentment and amplifies fear-inducing elements in the creepy house. Neural Style Transfer heavily relies on randomly selected reference images. While it produces artistic textures aligned with target emotions, such as Row 1, Column 3, it lacks in preserving essential semantic content in the source images. CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Rows 1-3, Column 4. Ip2p demonstrates limited image generation abilities with single-word emotion prompts. For example, the Halloween scene in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 2, Column 5). Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions. This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 6, Column 6.
In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible. For instance, in Row 1, Column 7, our method ingeniously transforms an unpleasant lunch box into a flower. In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
@@ 5.3 Generalization to Real-world Scenarios
Editing images within the same emotional valence presents greater challenges compared to cross-valence emotion image editing due to subtle differences in emotional cues. Figure 7(a) showcases our EmoEditor’s results within the same valence. Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes. In Example 1, it transitions from awe to contentment, depicting a sunset scene with afterglow. In Example 2, it generates a blooming white flower with additional purple flowers to evoke excitement. In Example 3, transitioning from amusement to awe, it transforms a lively spring scene into a tranquil night scene with moonlight.
In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all. We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset [16], commonly used for object detection. These images depict everyday life scenes with neutral emotions. Our EmoEditor applies positive and negative emotion transformations to these source images. Figure 7(b) illustrates the generated results. In Example 1, to inject amusement into an office space, our EmoEditor embellishes the walls with delightful watercolor paintings and infuses lively and vibrant colors into the room furniture. Conversely, to evoke fear, our model creates an aged desk and dimly lit walls. In Example 2, for an image featuring a courtyard with a teddy bear observing the garden, our method enriches the scene by adding a beautiful flower bed, creating an awe-inspiring effect. Conversely, to evoke feelings of disgust, our EmoEditor removes the flowers from the courtyard and transforms the view into a dilapidated scene with shabby clothes hanging on the walls and trash bags in the corner. It is crucial to emphasize that our EmoEditor is not trained in these neutral scenarios. All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
@@ 6 Discussion
We present a new and important problem of emotion-evoked image generation. To address this, we present EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics. Moreover, we contribute EmoPair dataset for model training. To benchmark all the methods, we introduce new evaluation metrics and establish standard protocols in emotion understanding, visualization, and reasoning. While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images. See Supp. for details.
@@
In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2p[3] and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 [4]. Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten. Figure S1 illustrates examples of these instructions. To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
@@
S2.1 Baseline: Large Model Series
Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS). This includes BLIP [15] for image captioning, followed by GPT-3 [4] for text instruction generation, and Ip2p for image editing based on the instructions.
@@
S2.2 Human Psychophysics Experiment
@@
We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)[26] (Online). We recruited 417 participants, each undergoing 100 trials. Additionally, we conducted in-lab experiments, involving 10 volunteers. Similarly, each participant reviewed 100 trials. The consistency between online and in-lab experimental results demonstrates the superiority of our approach in producing results that evoke the target emotion compared to other state-of-the-art methods.
Figure S6: Preference of Our Method Over SOTA Methods. We conducted both online experiments (417 participants) and in-lab experiments (10 participants). The left column represents the online results, while the right column displays the in-lab results. The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions. Chance is 50% in the red dotted line.
@@
Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-Transfer[7] (NST) and CLIP-Styler[14] (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes. Additionally, our EmoEditor surpasses Color-transfer[21] (CT) and Ip2p[3] in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions. This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones. Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S7 and S8. The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories. From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer [7] and CLIP-Styler [14] more in the manipulation of emotions from positive to negative. This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image. In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images compared to these two approaches.

### v2

@@ 4 Experiment
We selected 504 images randomly from the EmoSet dataset Yang et al. (2023), ensuring they were distinct from those in our EmoPair dataset. According to the emotion labels of EmoSet, there are 63 images for each source emotion category. Subsequently, we evaluated both competitive baselines and our EmoEditor on these source images and presented the results in the following subsections.
@@
Ideally, with source images and their classified emotions, our goal is to generate images evoking the other 7 target emotions. Considering emotional valence, we divided the test set into two subsets: within-valence emotion transfer and cross-valence emotion transfer. This division serves two purposes. Firstly, cross-valence emotion transfer is more common in real-world scenarios, like enhancing positive emotions in stressful environments or creating fear in lively settings during Halloween. Secondly, images of the same valence may contain subtle differences challenging for humans to discern. For instance, an amusement park scene might evoke a mixture of excitement and amusement.
@@
Baselines. We compare our EmoEditor with five state-of-the-art methods: (1) Color-transferPitié et al. (2007) (CT); (2) Neural-Style-TransferGatys et al. (2015) (NST). Both methods require an additional reference image from the target emotion category as inputs. We randomly selected these reference images from the EmoSet dataset. We also include representative text-to-image editing models (3) CLIP-StylerKwon, Ye (2022) (Csty) and (4) Ip2p Brooks et al. (2023). We provide these models with the target emotion category as a one-word text prompt, such as “excitement". Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline (5) “Large Model Series" (LMS). This includes BLIP Li et al. (2022) for image captioning, followed by GPT-3 Brown et al. (2020) for text instruction generation, and Ip2p for image editing based on the instructions (See Supp.Sec.S2.3).
@@
[Uncaptioned image] Figure 5: Human Psychophysics Experiment Results. The average proportion of images that human participants prefer our EmoEditor over other methods is 56%. Chance is 50% (red dotted line). Error bars are standard errors.    Method EMR(%)↑ ESR(%)↑ ENRD↓ ESS↓ CT 6.89 79.32 33.29 7.36 NST 34.42 92.01 34.45 18.57 Csty 11.51 85.52 41.47 36.64 Ip2p 2.53 67.76 9.39 12.71 LMS 11.51 77.38 26.13 19.74 w/o $\mathcal{P}$ 5.06 69.15 19.45 14.93 w/o $L_{emb}$ 43.35 91.62 22.53 16.00 Ours 50.20 92.86 23.98 16.27 Table 1: Quantitative Evaluation of Generated Images for All Competitive Methods and Ablated Models of our EmoEditor. See Sec.4 for the methods and evaluation metrics. See Sec.5.1 for the result analysis.
Human Psychophysics Experiment. We evaluate the generated results of all methods on Amazon Mechanical Turk (MTurk)Turk (2012). We recruit 417 participants, with each participant undergoing 100 trials, yielding total 417,000 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. All participants are properly compensated. In each trial, participants engage in a “two-alternative forced choice" task. They are presented with two image stimuli and must select the one that most strongly evokes the target emotion in them. The two image stimuli consists of a result generated by our EmoEditor and a randomly sampled result from the five baselines introduced above. The trial presentation order and the binary choices for each trial are randomized. To assess the quality of data collection, we introduce control trials and use them as filtering criteria. We discard the data from those people who fail the control trials at least 17% of the time. See Supp.Sec.S2.4 for more details.
Evaluation Metrics. While CLIP Radford et al. (2021) is commonly used to assess text-image consistency, it doesn’t fully capture the complexity of emotions Widhoelzl, Takmaz (2024). Relying on CLIP for emotional alignment may overlook important nuances in images. Thus, we introduce four novel evaluation metrics to quantitatively assess emotion-evoked image generation performance.
Emotion Matching Ratio (EMR) and Emotion Similarity Ratio (ESR) assess the extent to which the generated images evoke target emotions. EMR: Similar to Goetschalckx et al. (2019), we also use our emotion predictor $\mathcal{P}$ to assess the emotional category of generated images and determine their alignment with target emotions. We acknowledge that using $\mathcal{P}$ as an evaluator may introduce favorable biases toward our model. However, the absence of existing emotion evaluators leaves us no choice but to rely on $\mathcal{P}$. ESR: $\mathcal{P}$ generates emotion probability distributions $e_{g}$ and $e_{s}$ for generated and source images. We compute ESR using Kullback-Leibler Divergence distance $KLD(\cdot)$, where $KLD(e_{g},e_{oh})<KLD(e_{s},e_{oh})$. $e_{oh}$ is the binary one-hot target emotion vector, indicating the impact of edits on generated images transitioning from the source emotion to the target emotion.
Emotion-Neutral Region Deviation (ENRD) and Edge Structure Similarity (ESS) assess the structural coherence and semantic consistency between source and generated images. ENRD: Using Grad-CAM Selvaraju et al. (2017) with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify emotionally neutral regions on source images, valued at 0. Then, we compute the pixel-level L1 distance between these regions on source and generated images. ESS: Employing the Canny edge detection algorithm with thresholds of 200 and 500 on both source and generated images, we determine the L1 norm between their edge maps to quantify structural disparities.
@@ 5.1 Quantitative Evaluation in Cross-Valence Scenarios
Figure 5 shows human psychophysics experimental results (see Supp.Sec.S3.1 for more results). If all methods generate images that elicit target emotions equally effectively, human participants would make random choices, with 50% chance in the “two-alternative forced choice" tasks. Complimentary to psychophysics experiments, Table 1 presents quantitative results using four metrics from Sec.4. Observing Figure 5, we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines, indicating its proficiency in evoking target emotions. This is supported by EmoEditor’s highest EMR and ESR scores in Table 1. Additionally, EmoEditor achieves moderately lower ENRD and ESS, suggesting maintained structural and semantic coherence in the generated images.
@@
Figure 6: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green. See Supp.Fig.S10 for more examples.
From Figure 5, we noted that our EmoEditor significantly outperforms Color Transfer (CT), indicating that emotion-evoked image generation involves more than global color and brightness adjustments. CT also has lower EMR and ESR than our EmoEditor in Table 1. As expected, CT only transfers colors, preserving the edges of the source image and hence, resulting in the lowest ESS scores. Similarly, our EmoEditor is significantly favored over Ip2p by human participants, indicating that current text-to-image diffusion models face challenges in understanding emotional cues and generating effective text instructions for editing emotion-evoked images. Despite Ip2p having comparable ENRD and ESS scores to our EmoEditor, its significantly lower EMR and ESR scores suggest that it makes minimal changes to source images, maintaining high structural and semantic similarity but failing to evoke target emotions. Additionally, our EmoEditor also outperforms Large Model Series (LMS) based on human preference scores in Figure 5, 39% in EMR, 15% in ESR, 2 in ENRD, and 3 in ESS in Table 1. This suggests that LMS accumulates biases and errors across all procedural steps, highlighting the importance of end-to-end training for emotion-evoked image generation.
Neural-Style-Transfer (NST) and CLIP-Styler (Csty) excel in inducing negative emotions through distortion and irregular textures on source images. They are preferred in positive-to-negative emotion manipulations, primarily in the upper quadrants of their confusion matrix (see Supp.Fig.S8). However, these methods render images challenging to comprehend, indicated by higher ENRD and ESS scores in Table 1, suggesting the distorted images lose structural similarities and semantic content. In contrast, our EmoEditor outperforms both methods in generating positively valenced images, with higher preference scores in the lower quadrants of the confusion matrix (see Supp.Fig.S8).
In ablation studies (Table 1), removing either the emotion predictor $\mathcal{P}$ (Row 6) or the alignment loss $\lambda L_{emb}$ (Row 7) significantly impairs performance. This leads to lower EMR and ESR scores, highlighting the crucial roles of both components in our EmoEditor for detecting emotional cues from source images, emulating human-like reasoning in generating editing instructions, and executing iterative refinements during inference. We also perform ablation studies on different max iterations for iterative emotion inference, showing that this iterative process improves the model’s performance, achieving stable and high-quality emotionally evocative images (see Supp.Sec.S3.3).
@@ 5.2 Visualization of Emotional Image Generation across Valence
Figure 1 and 6 present visualizations of generated images for all competitive methods in cross-valence scenarios (see Supp.Sec.S3.2 for more results). Color Transfer primarily replicates tonal characteristics from a reference image, often ineffective for significant emotional augmentation. Neural Style Transfer heavily relies on randomly selected reference images. While it produces artistic textures aligned with target emotions, it lacks in preserving essential semantic content in the source images. CLIPstyler excels in producing negative emotions but introduces challenging-to-interpret textures, as seen in Column 4. Ip2p demonstrates limited image generation abilities with single-word emotion prompts. For example, the fire in the source image remains almost identical to the generated image, due to the inability to comprehend emotions (Row 1, Column 5). Even after concatenating a series of large vision and language models before Ip2p, LMS does not always generate desirable outputs, triggering target emotions. This can be seen by the generated image of a smiling face, failing to evoke sadness in Row 2, Column 6.
In contrast, our EmoEditor needs only the target emotion and source image inputs, yet produces highly creative images that evoke target emotions while striving to maintain scene structures and semantic coherence as much as possible. For instance, in Row 1, Column 7, our method does not change the subject of the fire but turns the fire in the woods into a bonfire and generates a beautiful night sky in the background, thus evoking contentment. In Row 6, Column 7, it accurately adjusts the character’s facial expression while retaining the original content elements.
@@ 5.3 Generalization to Real-world Scenarios
Editing images within the same emotional valence presents greater challenges compared to cross-valence emotion image editing due to subtle differences in emotional cues. Figure 7(a) showcases our EmoEditor’s results within the same valence. Our EmoEditor displays a nuanced understanding of emotional cues and produces impressive creative outcomes. Example 1 transitions from awe to contentment, depicting a sunset scene with afterglow. Example 2 generates a blooming white flower with additional purple flowers to evoke excitement. Example 3 transforms a lively spring scene into a tranquil night scene with moonlight to evoke awe.
In real-world scenarios, some naturalistic images may evoke mixed emotions or no strong emotional responses at all. We challenge our EmoEditor to generate emotion-evoked images using neutral-valence source images randomly selected from the MSCOCO dataset Lin et al. (2014), commonly used for object detection. These images depict everyday life scenes with neutral emotions. Our EmoEditor applies positive and negative emotion transformations to these source images. Figure 7(b) illustrates the generated results. In Example 1, to inject amusement into an office space, our EmoEditor embellishes the walls with delightful watercolor paintings and infuses lively and vibrant colors into the room furniture. Conversely, to evoke fear, our model creates an aged desk and dimly lit walls. In Example 2, for an image featuring a courtyard with a teddy bear observing the garden, our method enriches the scene by adding a beautiful flower bed, creating an awe-inspiring effect. Conversely, to evoke feelings of disgust, our EmoEditor removes the flowers from the courtyard and transforms the view into a dilapidated scene with shabby clothes hanging on the walls and trash bags in the corner. It is crucial to emphasize that our EmoEditor is not trained in these neutral scenarios. All the generated results stem solely from the model’s own creativity and imagination, without any human-annotated editing instructions or scene interpretations.
@@ 6 Discussion
We present a new and important problem of emotion-evoked image generation. To address this, we present EmoEditor, an image diffusion model that understands emotional cues, creates implicit editing instructions aligned with human decisions, and manipulates image regions to evoke emotions, while maintaining coherent scene structures and semantics. Moreover, we contribute EmoPair dataset for model training. To benchmark all the methods, we introduce new evaluation metrics and establish standard protocols in emotion understanding, visualization, and reasoning. While our EmoEditor demonstrates superior quantitative performance and visually striking image generation results, we acknowledge several limitations in accurately handling fine-grained details of visual features on small faces within crowded scenes and generating emotion-evoked images without exacerbating semantic and structural disparities between source and target images. See Supp.Sec.S3.4 for details.
@@
In response to the absence of an image dataset specifically designed for emotion-evoke image generation, we introduce EmoPair, comprising two distinct subsets: the EmoPair-Annotated Subset (EPAS), encompassing 331,595 image pairs sourced from Ip2pBrooks et al. (2023) and annotated with emotion labels; and the EmoPair-Generated Subset (EPGS), featuring 6,949 pairs generated through text instructions specifying target emotions.
For the EmoPair-Generated Subset (EPGS), we formulated 50 general instructions that are agnostic to source images, prompting transitions to desired emotions across 8 categories using GPT-3 Brown et al. (2020). Human annotators then ranked these instructions based on efficacy within each emotion category, ultimately retaining the top ten. Figure S1 illustrates examples of these instructions.
We employ the following selection criteria to control the quality of generated image pairs of EPGS: (1) Using our emotion predictor $\mathcal{P}$, we analyze the generated images and only select those with a Top-1 classification confidence over 90% for the target emotions. (2) To ensure the preservation of similar scene structures, we utilize Structural Similarity Index (SSIM) Wang et al. (2004) and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) for filtering the remaining outcomes. SSIM measures structural similarity between images, while LPIPS quantifies perceptual differences. Specifically, we require the generated images $\hat{x}$ and the source image $x$ to meet the conditions below: 0.3 < SSIM($x$, $\hat{x}$) < 0.6 and LPIPS($x$, $\hat{x}$) > 0.1. Ultimately, EPGS retains 6,949 image pairs.
To provide a more comprehensive overview of our dataset, additional image pair examples are presented in Figure S2 and S3.
@@
S2.1 Emotion Encoder
Our Emotion Encoder $\tau_{\theta}$ is a fully connected network designed for transforming one-hot encoded emotion input vectors into the structured emotion embedding $e$. The architecture of the model is composed of a series of fully connected layers that progressively increase the dimensionality of the input. Figure S4 shows the network structure of our emotion encoder $\tau_{\theta}$.
The network begins with an input layer of size 8, which is then passed through a sequence of fully connected layers with increasing sizes: 256, 512, and 768 neurons, respectively. Each of these layers is followed by a ReLU activation function to introduce non-linearity and improve the model’s learning capacity. The final linear layer projects the output to a dimension of $77\times 768$, effectively structuring the embedding to fit the input size of our denoising network $\epsilon_{\theta}$.
Figure S4: The network structure of our emotion encoder $\tau_{\theta}$.
S2.2 Implementation Details
Follow Wang et al. (2022), we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p Brooks et al. (2023) weights. Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$. The frozen emotion predictor $\mathcal{P}$ is only used during inference. We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer. The max time step $T$ is set to 1,000.
S2.3 Baseline: Large Model Series
Emotion-evoked image generation involves three key steps: image understanding, instruction generation, and instruction-based image editing. To tackle these, we concatenate existing large language and vision models and introduce the baseline “Large Model Series" (LMS). This includes BLIP Li et al. (2022) for image captioning, followed by GPT-3 Brown et al. (2020) for text instruction generation, and Ip2p for image editing based on the instructions.
@@
S2.4 Human Psychophysics Experiment
@@
We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)Turk (2012) (Online). We recruited 417 participants, each undergoing 100 trials. Additionally, we conducted in-lab experiments, involving 10 volunteers. Similarly, each participant reviewed 100 trials. The consistency between online and in-lab experimental results demonstrates the superiority of our approach in producing results that evoke the target emotion compared to other state-of-the-art methods.
Figure S7: Preference of Our Method Over SOTA Methods. We conducted both online experiments (417 participants) and in-lab experiments (10 participants). The left column represents the online results, while the right column displays the in-lab results. The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions. Chance is 50% in the red dotted line. Error bars are standard errors.
@@
Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-TransferGatys et al. (2015) (NST) and CLIP-StylerKwon, Ye (2022) (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes. Additionally, our EmoEditor surpasses Color-transferPitié et al. (2007) (CT) and Ip2pBrooks et al. (2023) in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions. This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones. Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S8 and S9. The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories. From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer Gatys et al. (2015) and CLIP-Styler Kwon, Ye (2022) more in the manipulation of emotions from positive to negative. This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image. In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images compared to these two approaches.